OpenStudy (anonymous):

What's the difference between linear and exponential functions? @poojah can you help?

OpenStudy (decarr432):

Hello!

OpenStudy (anonymous):

Hi @Decarr432

OpenStudy (decarr432):

Exponential means a function whose value is a constant raised to the power of the argument

OpenStudy (decarr432):

Linear functions are those whose graph is a straight line

OpenStudy (xapproachesinfinity):

so this function \(x^x \) is not exponential

OpenStudy (anonymous):

So what is x^x? is it linear?

OpenStudy (xapproachesinfinity):

no it is not linear

OpenStudy (solomonzelman):

Exponential function is in the form of: \(\color{#000000}{\displaystyle y=a(b)^x}\) And the linear function is in the form of: \(\color{#000000}{\displaystyle y=mx+b}\) (m could be 0, too) Ok, but still, what is the difference? Let me give you a hint for this.s s The table for: \(\color{#000000}{\displaystyle y=a(b)^x}\) \(\color{#000000}{\displaystyle x}\) \(\color{#000000}{\displaystyle y}\) \(\color{#000000}{\displaystyle 1}\) \(\color{#000000}{\displaystyle a\cdot b }\) \(\color{#000000}{\displaystyle 2}\) \(\color{#000000}{\displaystyle a\cdot b^2 }\) \(\color{#000000}{\displaystyle 3}\) \(\color{#000000}{\displaystyle a\cdot b^3 }\) \(\color{#000000}{\displaystyle 4}\) \(\color{#000000}{\displaystyle a\cdot b^4 }\) (So wouldn't I multiply times b every time x goes up by 1?) The table for: \(\color{#000000}{\displaystyle y=mx+b}\) \(\color{#000000}{\displaystyle x}\) \(\color{#000000}{\displaystyle y}\) \(\color{#000000}{\displaystyle 1}\) \(\color{#000000}{\displaystyle m+b }\) \(\color{#000000}{\displaystyle 2}\) \(\color{#000000}{\displaystyle 2m+b }\) \(\color{#000000}{\displaystyle 3}\) \(\color{#000000}{\displaystyle 3m+b }\) \(\color{#000000}{\displaystyle 4}\) \(\color{#000000}{\displaystyle 4m+b }\) (So wouldn't I add an "m" every time x goes up by 1?)